6. Evaluate the surface integral // F.ds where the surface S is the sphere x2 + y2 + z2 = 4 [ Ꭻ Ꭻs . and F = (xz, -2y, 3.c) w7. Use the Divergence Theorem to recalculate the surface integral in problem 6.
Answer #1

Please see the attched image for solution

6 = <nz, -ey, 3n) 7= 2 singus di + 2512051n09 + Plosok x² + y² + 2 = 4. = 2² 2. asinouso 4 26100 6100% 7. 2660 SSF do = SS El

ery of sido coso cosd dodø - cost su 20 dodo he LOX 20 = 1-281920 2. 2. = cosbx (0 - 59640)5 dd if= -60520 se soodlot = 0 8

& SSF. dr = SSS OF do DE 2(22) + 2(-2y) +2/32) -2. =rausso f 20000 2 hool D. F rioso - 2. ss of dr = * 56050-2) ** sino doda